OpenStudy (anonymous):

In isosceles trapezoid JKLM, Leg JK = 5x + 10, base KL = 6x + 2, and leg LM = 2x +8. Find the value of x.

OpenStudy (anonymous):

check this one: http://mathworld.wolfram.com/IsoscelesTrapezoid.html because trapezoid isosceles - it's means that: 5x+10=2x+8 solve it for x

OpenStudy (anonymous):

3x=-2 x=-2/3

OpenStudy (anonymous):

JK and LM are equal 5x + 10 = 2x + 8 subtract 10 from both sides 5x = 2x -2 Subtract 2x from both sides 3x = -2 Divide both by 3 x = -2/3